how do you solve the inequality, write the solution set in interval notation then graph it? Example: 24n+20 is< or = to 4(5n+13)
Answer by azn_hannah(3)


Hello, my name is Hannah and I would like to help you solve the inequality: 24n+20 is < or = 4(5n+13)
First of all, you are trying to isolate n, keep that in mind. You must simplify the equation by using the distibutive property.
24n+20 is < or = 20n+52
Next, because the equation has varibles on both sides, you must collect the variables to form one. In this case, you would subtract 20n from both sides.
24n+20-20n is < or = 20n+52-20n
You are left with the equation 4n+20 is < or = 52. In order to isolate n, you have to take out additional information (addition, subtraction, etc.).
4n+20-20 is < or = 52-20
Now you have 4n is < or = 32 as your equation. Now, divide each side of the equation by 4.
4n/4=32/4
The answer is x is < or = 8.
If you need anything else, please inform me. I'll be more than happy to help. =]